CRY+VIS Interface Meeting
Date: 2021/7/9 13:00-14:00
Participants: Takahashi, Ikeda, Sato, Ushiba, Kimura, Aso, Tomaru, Node, Hirata
- IX: To the finalization.
- EX: Alignment for the green beam was finished. To the finalization
- IY: Height adjustment is on going.
- EY: Mirror cleaning is on going in Mitaka. Measurement system had a problem. The shipping will be in the end of next week at earliest.
- Corrugate tube fixing jig: Implementation is scheduled on 7/12-13.
- TM height setting: ITMY.
- Commissioning in air: X-arm.
- Commissioning in vacuum:
- Overall commissioning:
- Acceptance check: Every Friday.
- PF actuator: After SC.
- Separation of rack in 2F: This is scheduled on 7/16.
- New vacuum pumps for V-chambers: Considering the port location.
- Extension of cables for picomotor drivers: Tested in EYV.
Adjustable height range of the filter chain in IY
- The condition of F1 stepper motor is the same as the F3 in IX. We tried to move the motor with the stand alone driver, but the FR is still not working.
- The motor is still not working?
- The FR slider is stuck?
- The range of BF is too small. Check the mechanical stopper.
- Tune the F0 more to lower frequency (now 0.4Hz) in order to extend the range.